# Phase transitions for unique codings of fat Sierpinski gasket with multiple digits

### v1

@@ 2. Characterization of $\widetilde{U}_{\beta,M}$
In view of Figure 4 (left graph), we can also prove (2.5).
@@
In view of Figure 4 (middle graph), this yields (2.5).
@@
In view of Figure 4 (right graph), this also proves (2.5). ∎
@@
Suppose $\beta\in(1,\frac{M+2}{2}]$. Then $S_{\beta,M}=\Delta_{\beta,M}$, and in this case each point in the overlap region $O_{\beta,M}$ has at least two codings (cf. [23, Proposition 2.4]). So, by (1.2) and (1.3) it follows that $\widetilde{U}_{\beta,M}=U_{\beta,M}$, and also $\widetilde{{\mathbf{U}}}_{\beta,M}={\mathbf{U}}_{\beta,M}$ since $\Pi_{\beta}$ is bijective. ∎
@@ 3. Generalized golden ratio $\beta_{G}(M)$ for a general $M$
The following result shows that $\beta_{G}(M)$ is the first critical base for $\widetilde{{\mathbf{U}}}_{\beta,M}$, i.e., $\widetilde{{\mathbf{U}}}_{\beta,M}$ is finite if and only if $\beta\leq\beta_{G}(M)$. When $M=1$, this was proven by Sidorov [28, Theorem 4.1] (see also, [23]).
@@
#$\#\widetilde{{\mathbf{U}}}_{\beta,M}<+\infty\quad\Longleftrightarrow\quad\beta\leq\beta_{G}(M).$
@@
Case (I). $M=3N+1$ with $N\in\mathbb{N}_{0}$. Then by (3.1), Lemma 2.1 and Proposition 2.3 we only need to prove that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{M0}^{\infty},\alpha_{0M}^{\infty}\right\}$ if $\delta(\beta)\preccurlyeq((N+1)NN)^{\infty}$; and #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$ if $\delta(\beta)\succ((N+1)NN)^{\infty}$.
Take $\beta\in(1,M+1]$ so that $\delta(\beta)\preccurlyeq((N+1)NN)^{\infty}$. Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$, and suppose $d_{1}^{1},d_{1}^{2},\overline{d_{1}^{\oplus}}<M$. Then by Proposition 2.3 and (2.1) it follows that
@@
which yields $\overline{d_{2}^{\oplus}d_{3}^{\oplus}d_{4}^{\oplus}}=(N-1)(N+1)(N+1)$. This leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. For (3.3) we assume by (2.7) that $d_{2}=\alpha_{(N+1)N}$. Then by Proposition 2.3 and (2.1) it follows that
@@
If $d_{3}=\alpha_{(N-1)(N+1)}$, then by a similar argument as in (3.2) we will obtain that $d_{3}d_{4}d_{5}=\alpha_{(N-1)(N+1)}\alpha_{(N+1)N}\alpha_{(N+1)N}$, leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. So, $d_{2}d_{3}\in\left\{\alpha_{(N+1)N}\alpha_{N(N+1)},\alpha_{(N+1)N}\alpha_{NN}\right\}$. By Proposition 2.3 and the same argument as above we can deduce that
@@
which again leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$.
@@
So #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$.
Case (II). $M=3N+2$ with $N\in\mathbb{N}_{0}$. By (3.1), Lemma 2.1 and Proposition 2.3 it suffices to prove that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{M0}^{\infty},\alpha_{0M}^{\infty}\right\}$ if $\delta(\beta)\preccurlyeq((N+1)(N+1)N)^{\infty}$; and #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$ if $\delta(\beta)\succ((N+1)(N+1)N)^{\infty}$.
@@
Both cases will lead to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. Similar to Case $M=3N+1$, we can deduce that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{M0}^{\infty},\alpha_{0M}^{\infty}\right\}$. On the other hand, if $\beta\in(1,M+1]$ with $\delta(\beta)\succ((N+1)(N+1)N)^{\infty}$, then we can prove that
@@
So #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$.
Case (III). $M=3N+3$ with $N\in\mathbb{N}_{0}$. By (3.1), Lemma 2.1 and Proposition 2.3 it suffices to prove that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{0M}^{\infty},\alpha_{M0}^{\infty}\right\}$ if $\delta(\beta)\preccurlyeq(N+1)^{\infty}$, and #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$ if $\delta(\beta)\succ(N+1)^{\infty}$.
@@
leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. So, $d_{1}^{1}=M,d_{1}^{2}=M$ or $\overline{d_{1}^{\oplus}}=M$. By the same argument as that for $M=3N+1$ we can prove $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{0M}^{\infty},\alpha_{M0}^{\infty}\right\}$. On the other hand, take $\beta\in(1,M+1]$ so that $\delta(\beta)\succ(N+1)^{\infty}$. Then one can verify by Proposition 2.3 that
@@
which implies #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$. This completes the proof. ∎
@@ 4. Generalized Komornik-Loreti constant $\beta_{c}(M)$ for $M=3N+1$
Let $M=3N+1$ with $N\in\mathbb{N}_{0}$. Recall from Definition 1.2 the generalized Komornik-Loreti constant $\beta_{c}(M)$. In this section we will show that $\dim_{H}\widetilde{{\mathbf{U}}}_{\beta,M}>0$ if and only if $\beta>\beta_{c}(M)$; and prove Theorem 1.4 for $M=3N+1$. When $M=1$, Theorem 1.4 was established by Kong and Li [23]. In this section we may assume $M=3N+1$ with $N\in\mathbb{N}$. The new challenge is that for $M=3N+1$ with $N\geq 1$ the digit $N-1$ might occur in $d_{n}^{1},d_{n}^{2}$ or $\overline{d_{n}^{\oplus}}$ for a sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$, while for $M=1$ this can never happen. Fortunately, we show in Lemma 4.5 that this digit $N-1$ can occur at most 3 times in any sequence of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$.
@@ 4.1. Properties of $\beta_{c}(M)$
$(\lambda_{i})_{i=1}^{\infty}=\sigma(\varphi_{1}(\tau_{0})\varphi_{1}(\tau_{1})\varphi_{1}(\tau_{2})\ldots),$
[…]
@@
[…]
$(\tau_{i})_{i=0}^{\infty}=\phi(\tau_{0})\phi(\tau_{1})\phi(\tau_{2})\ldots.$
[…]
$(\lambda_{i})_{i=0}^{\infty}=\varphi_{1}(\tau_{0})\varphi_{1}(\tau_{1})\varphi_{1}(\tau_{2})\ldots,$
@@
Next we show that $(\lambda_{i})$ is the quasi-greedy expansion of $1$ in base $\beta_{c}(M)$.
@@
Then for a sequence $\mathbf{a}=a_{1}a_{2}\ldots\in{\Sigma}_{1}^{\mathbb{N}}$ we set $\phi_{1}(\mathbf{a}):=\phi_{1}(a_{1})\phi_{1}(a_{2})\cdots\in\Sigma_{1}^{\mathbb{N}}$. We emphasize that each digit $a_{i}$ is a block of length $3$ from ${\Sigma}_{1}$. Note that $(\lambda_{i})\in\Sigma_{1}^{\mathbb{N}}$. We define its conjugacy by
@@
Since the proof of (4.7) is similar to (4.6), we only prove (4.6). Take $n\in\mathbb{N}_{0}$ and $i\in\left\{0,1,\ldots,3\cdot 2^{n}-1\right\}$. We will prove (4.6) by the following three cases: (I) $i=3k$, (II) $i=3k+1$, and (III) $i=3k+2$ for some $0\leq k<2^{n}$.
@@
Next we assume $10\leq k<2^{n}-$ 1. Then by (4.1) we have $\lambda_{i+1}=\lambda_{3k+2}=N<N+1=\lambda_{1}$, which yields $\lambda_{i+1}\ldots\lambda_{3\cdot 2^{n}}\prec\lambda_{1}\ldots\lambda_{3\cdot 2^{n}-i}$. On the other hand, observe by (4.1) that
@@
Case (III). $i=3k+2$ with $0\leq k<2^{n}$. If $k=2^{n}-1$, then $\lambda_{i+1}=\lambda_{3\cdot 2^{n}}=N+1$. Since $\lambda_{1}=N+1$ and $\hat{\lambda}_{1}=N$, it is clear that (4.6) holds for $i=3\cdot(2^{n}-1)+2$. In the following we assume $10\leq k<2^{n}-$ 1. Then by (4.1) we obtain
@@
Note by Definition 1.2 and (4.1) that $(\lambda_{i})$ is an expansion of $1$ in base $\beta_{c}=\beta_{c}(M)$. Then by Lemmas 2.1 and 4.3 it follows that $\delta(\beta_{c})=(\lambda_{i})$. So, by (4.1) we obtain
@@
Suppose on the contrary that $\beta_{c}\in(N+1,N+2)$ is algebraic. Then so is $\beta_{c}^{-3}\in(0,1)$. By Lemma 4.4 it follows that the left-hand side of (4.10) is transcendental, while the right hand-side of (4.10) is algebraic, leading to a contradiction. So, $\beta_{c}$ is transcendental. ∎
@@ 4.2. Patterns in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$
In the following we describe the block patterns appearing in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. Let
$\pazocal A_{1}$ :=$:=\left\{\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{(N+1)N},\quad\alpha_{NN}\alpha_{N(N+1)}\alpha_{NN},\right.$
@@
Comparing the definitions of $({\mathbf{t}}_{n})$ and $(\mathbf{a}_{n})$ in (4.11) and (4.14) respectively, it follows that
@@
for any $n\geq 0$.
The big difference between $M=1$ and $M=3N+1$ with $N\geq 1$ is that in the later case the digit $N-1$ might occur in $d_{n}^{1}=N-1,d_{n}^{2}=N-1$ or $\overline{d_{n}^{\oplus}}=N-1$ for some sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. In other words, when $M=3N+1$ with $N\geq 1$, sequences in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ might contain digit from
@@
#$\#\left\{i\in\mathbb{N}:d_{i}\in\mathscr{E}_{1}\right\}\leq 3.$
Proof.
@@
(4.17) $d_{m+k}^{1},\leavevmode\nobreak\ d_{m+k}^{2},\leavevmode\nobreak\ \overline{d_{m+k}^{\oplus}}\leq N+1<M\quad\forall k\geq 0.$
@@
For $d_{m+p+2}$ we have three possibilities: $d_{m+p+2}=\alpha_{(N+1)N},\alpha_{N(N+1)}$ or $\alpha_{(N+1)(N+1)}$.
@@
Suppose on the contrary that $d_{m+p+2}=\alpha_{(N+1)(N+1)}$. Then by using $d_{m+p}d_{m+p+1}d_{m+p+2}=\alpha_{(N+1)(N+1)}\alpha_{NN}\alpha_{(N+1)(N+1)}$, Propositions 2.3 and 4.2 we must have
@@
This implies that $\overline{d_{m+p+2}^{\oplus}}=N-1<M$ and $\overline{d_{m+p+3}^{\oplus}d_{m+p+4}^{\oplus}}=(N+1)(N+1)\succ\delta_{1}(\beta_{c})\delta_{2}(\beta_{c})$, leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, Claim 1 holds.
By Claim 1 we have $d_{m+p+2}=\alpha_{(N+1)N}$ or $\alpha_{N(N+1)}$. By symmetry we may assume $d_{m+p+2}=\alpha_{(N+1)N}$. Then by (4.17), Propositions 2.3 and 4.2 it follows that

### v2

@@ 2. Characterization of $\widetilde{U}_{\beta,M}$
In view of Figure 4a , we can also prove (2.5).
@@
In view of Figure 4b, this yields (2.5).
@@
In view of Figure 4c, this also proves (2.5). ∎
@@
Suppose $\beta\in(1,\frac{M+2}{2}]$. Then $S_{\beta,M}=\Delta_{\beta,M}$, and in this case each point in the overlap region $O_{\beta,M}$ has at least two codings (cf. [24, Proposition 2.4]). So, by (1.2) and (1.3) it follows that $\widetilde{U}_{\beta,M}=U_{\beta,M}$, and also $\widetilde{{\mathbf{U}}}_{\beta,M}={\mathbf{U}}_{\beta,M}$ since $\Pi_{\beta}$ is bijective. ∎
@@ 3. Generalized golden ratio $\beta_{G}(M)$ for a general $M$
The following result shows that $\beta_{G}(M)$ is the first critical base for $\widetilde{{\mathbf{U}}}_{\beta,M}$, i.e., $\widetilde{{\mathbf{U}}}_{\beta,M}$ is finite if and only if $\beta\leq\beta_{G}(M)$. When $M=1$, this was proven by Sidorov [29, Theorem 4.1] (see also, [24]).
@@
$\widetilde{{\mathbf{U}}}_{\beta,M}<+\infty\quad\Longleftrightarrow\quad\beta\leq\beta_{G}(M).$
@@
Case (I). $M=3N+1$ with $N\in\mathbb{N}_{0}$. Then by (3.1), Lemma 2.1 and Proposition 2.3 we only need to prove that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{M0}^{\infty},\alpha_{0M}^{\infty}\right\}$ if $\delta(\beta)\preccurlyeq((N+1)NN)^{\infty}$; and $\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$ if $\delta(\beta)\succ((N+1)NN)^{\infty}$.
Take $\beta\in(1,M+1]$ so that $\delta(\beta)\preccurlyeq((N+1)NN)^{\infty}$. Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. Recall that each $d_{i}=(d_{i}^{1},d_{i}^{2})$ and $\overline{d_{i}^{\oplus}}=M-d_{i}^{\oplus}$ with $d_{i}^{\oplus}=d_{i}^{1}+d_{i}^{2}$. Suppose $d_{1}^{1},d_{1}^{2},\overline{d_{1}^{\oplus}}<M$. Then by Proposition 2.3 and (2.1) it follows that
@@
which yields $\overline{d_{2}^{\oplus}d_{3}^{\oplus}d_{4}^{\oplus}}=(N-1)(N+1)(N+1)$. This leads to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. For (3.3) we assume by (2.7) that $d_{2}=\alpha_{(N+1)N}$. Then by Proposition 2.3 and (2.1) it follows that
@@
If $d_{3}=\alpha_{(N-1)(N+1)}$, then by a similar argument as in (3.2) we will obtain that $d_{3}d_{4}d_{5}=\alpha_{(N-1)(N+1)}\alpha_{(N+1)N}\alpha_{(N+1)N}$, leading to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. So,
$d_{2}d_{3}\in\left\{\alpha_{(N+1)N}\alpha_{N(N+1)},\alpha_{(N+1)N}\alpha_{NN}\right\}.$
By Proposition 2.3 and the same argument as above we can deduce that
@@
which again leads to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$.
@@
So $\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$.
Case (II). $M=3N+2$ with $N\in\mathbb{N}_{0}$. By (3.1), Lemma 2.1 and Proposition 2.3 it suffices to prove that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{M0}^{\infty},\alpha_{0M}^{\infty}\right\}$ if $\delta(\beta)\preccurlyeq((N+1)(N+1)N)^{\infty}$; and $\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$ if $\delta(\beta)\succ((N+1)(N+1)N)^{\infty}$.
@@
Both cases will lead to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. Similar to Case $M=3N+1$, we can deduce that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{M0}^{\infty},\alpha_{0M}^{\infty}\right\}$. On the other hand, if $\beta\in(1,M+1]$ with $\delta(\beta)\succ((N+1)(N+1)N)^{\infty}$, then we can prove that
@@
So $\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$.
Case (III). $M=3N+3$ with $N\in\mathbb{N}_{0}$. By (3.1), Lemma 2.1 and Proposition 2.3 it suffices to prove that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{0M}^{\infty},\alpha_{M0}^{\infty}\right\}$ if $\delta(\beta)\preccurlyeq(N+1)^{\infty}$, and $\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$ if $\delta(\beta)\succ(N+1)^{\infty}$.
@@
leading to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. So, $d_{1}^{1}=M,d_{1}^{2}=M$ or $\overline{d_{1}^{\oplus}}=M$. By the same argument as that for $M=3N+1$ we can prove $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{0M}^{\infty},\alpha_{M0}^{\infty}\right\}$. On the other hand, take $\beta\in(1,M+1]$ so that $\delta(\beta)\succ(N+1)^{\infty}$. Then one can verify by Proposition 2.3 that
@@
which implies $\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$. This completes the proof. ∎
@@ 4. Generalized Komornik-Loreti constant $\beta_{c}(M)$ for $M=3N+1$
Let $M=3N+1$ with $N\in\mathbb{N}_{0}$. Recall from Definition 1.2 the generalized Komornik-Loreti constant $\beta_{c}(M)$. In this section we will show that $\dim_{H}\widetilde{{\mathbf{U}}}_{\beta,M}>0$ if and only if $\beta>\beta_{c}(M)$; and prove Theorem 1.4 for $M=3N+1$. When $M=1$, Theorem 1.4 was established by Kong and Li [24]. In this section we may assume $M=3N+1$ with $N\in\mathbb{N}$. The new challenge is that for $M=3N+1$ with $N\geq 1$ the digit N1N-1 might occur in $d_{n}^{1},d_{n}^{2}$ or $\overline{d_{n}^{\oplus}}$ for a sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$, while for $M=1$ this can never happen. Remember that each digit $d_{i}=(d_{i}^{1},d_{i}^{2})$ and $\overline{d_{i}^{\oplus}}=M-d_{i}^{\oplus}$ with $d_{i}^{\oplus}=d_{i}^{1}+d_{i}^{2}$. Fortunately, we show in Lemma 4.5 that this digit N1N-1 can occur at most 3 times in any sequence of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$.
@@ 4.1. Properties of $\beta_{c}(M)$
$(\lambda_{i})_{i=1}^{\infty}=\sigma(\varphi_{1}(\tau_{0})\varphi_{1}(\tau_{1})\varphi_{1}(\tau_{2})\cdots),$
where $\sigma$ is the left-shift map, and $\varphi_{1}$ is the substitution defined by $0\mapsto N(N+1)N;\penalty 10000\ 1\mapsto(N+1)NN.$
@@
for all $k\geq 0$. Note that $(\tau_{i})_{i=0}^{\infty}$ is the fixed point of the substitution $\phi:0\mapsto 01;\penalty 10000\ 1\mapsto 10$ (cf. [3, Ch. 6]). That is
$(\tau_{i})_{i=0}^{\infty}=\phi(\tau_{0})\phi(\tau_{1})\phi(\tau_{2})\cdots.$
Therefore, comparing (4.2) with the definition of $\varphi_{1}:0\mapsto(0+N)(1+N)N;\penalty 10000\ 1\mapsto(1+N)(0+N)N$ we obtain that
$(\lambda_{i})_{i=0}^{\infty}=\varphi_{1}(\tau_{0})\varphi_{1}(\tau_{1})\varphi_{1}(\tau_{2})\cdots,$
@@
Next we show that $(\lambda_{i})$ is the quasi-greedy expansion of $1$ in the base $\beta_{c}(M)$.
@@
Then for a sequence $\mathbf{a}=a_{1}a_{2}\ldots\in{\Sigma}_{1}^{\mathbb{N}}$ we set $\phi_{1}(\mathbf{a}):=\phi_{1}(a_{1})\phi_{1}(a_{2})\cdots\in\Sigma_{1}^{\mathbb{N}}$. We emphasize that each digit $a_{i}$ is a block of length $3$ from ${\Sigma}_{1}$. Note that $(\lambda_{i})\in\Sigma_{1}^{\mathbb{N}}$. We define its conjugate by
@@
Since the proof of (4.7) is similar to that of (4.6), we only prove (4.6). Take $n\in\mathbb{N}_{0}$ and $i\in\left\{0,1,\ldots,3\cdot 2^{n}-1\right\}$. We will prove (4.6) by the following three cases: (I) $i=3k$, (II) $i=3k+1$, and (III) $i=3k+2$ for some $0\leq k<2^{n}$.
@@
Next we assume $0\leq k<2^{n}-1$. Then by (4.1) we have $\lambda_{i+1}=\lambda_{3k+2}=N<N+1=\lambda_{1}$, which yields $\lambda_{i+1}\ldots\lambda_{3\cdot 2^{n}}\prec\lambda_{1}\ldots\lambda_{3\cdot 2^{n}-i}$. On the other hand, observe by (4.1) that
@@
Case (III). $i=3k+2$ with $0\leq k<2^{n}$. If $k=2^{n}-1$, then $\lambda_{i+1}=\lambda_{3\cdot 2^{n}}=N+1$. Since $\lambda_{1}=N+1$ and $\hat{\lambda}_{1}=N$, it is clear that (4.6) holds for $i=3\cdot(2^{n}-1)+2$. In the following we assume $0\leq k<2^{n}-1$. Then by (4.1) we obtain
@@
Note by Definition 1.2 and (4.1) that $(\lambda_{i})$ is an expansion of $1$ in the base $\beta_{c}=\beta_{c}(M)$. Then by Lemmas 2.1 and 4.3 it follows that $\delta(\beta_{c})=(\lambda_{i})$. So, by (4.1) we obtain
@@
Suppose on the contrary that $\beta_{c}\in(N+1,N+2)$ is algebraic. Then so is $\beta_{c}^{-3}\in(0,1)$. By Lemma 4.4 it follows that the left-hand side of (4.10) is transcendental, while the right-hand side of (4.10) is algebraic, leading to a contradiction. So, $\beta_{c}$ is transcendental. ∎
@@ 4.2. Patterns in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$
Note that the blocks ${\mathbf{t}}_{n},n\geq 0$ are similar to the one-dimensional balanced words (see, e.g., [22, 23]). In terms of Proposition 2.3, we construct a sequence of two dimensional balanced words $(\mathbf{a}_{n})_{n\geq 0}$ appearing in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. These words $(\mathbf{a}_{n})$ are like basic bricks to describe the set $\widetilde{{\mathbf{U}}}_{\beta,M}$ for $\beta\leq\beta_{c}(M)$ with $M=3N+1$. Let
$\pazocal A_{1}$ $:=\left\{\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{(N+1)N},\quad\alpha_{NN}\alpha_{N(N+1)}\alpha_{NN},\right$.
@@
By comparing the definitions of $({\mathbf{t}}_{n})$ and $(\mathbf{a}_{n})$ in (4.11) and (4.14) respectively, it follows that
@@
for any $n\geq 0$, where for a word ${\mathbf{d}}=d_{1}\ldots d_{k}\in{\Omega}_{M}^{*}$ we write ${\mathbf{d}}^{1}=d_{1}^{1}\ldots d_{k}^{1}$, ${\mathbf{d}}^{2}=d_{1}^{2}\ldots d_{k}^{2}$ and $\overline{{\mathbf{d}}^{\oplus}}=\overline{d_{1}^{\oplus}}\ldots\overline{d_{k}^{\oplus}}$ with each $d_{i}=(d_{i}^{1},d_{i}^{2})$ and $\overline{d_{i}^{\oplus}}=M-d_{i}^{\oplus}=M-(d_{i}^{1}+d_{i}^{2})$.
The big difference between $M=1$ and $M=3N+1$ with $N\geq 1$ is that in the latter case the digit N1N-1 might occur in $d_{n}^{1}=N-1,d_{n}^{2}=N-1$ or $\overline{d_{n}^{\oplus}}=N-1$ for some sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. In other words, when $M=3N+1$ with $N\geq 1$, sequences in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ might contain digits from
@@
$\left\{i\in\mathbb{N}:d_{i}\in\mathscr{E}_{1}\right\}\leq 3.$
Before the proof we give an example to illustrate the main idea.
Example 4.6.
Let $M=4$. Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},4}\setminus\left\{\alpha_{40}^{\infty},\alpha_{04}^{\infty},\alpha_{00}^{\infty}\right\}$. Note that $\delta(\beta_{c})=212112111\ldots$. By Proposition 2.3 there exists $n_{0}\in\mathbb{N}$ such that for all $i\geq n_{0}$ we have either
$d_{i}\in\left\{\alpha_{21},\alpha_{12},\alpha_{11}\right\}\sim\left\{\left(\begin{array}[]{c}2\\ 1\\ 1\end{array}\right),\left(\begin{array}[]{c}1\\ 2\\ 1\end{array}\right),\left(\begin{array}[]{c}1\\ 1\\ 2\end{array}\right)\right\}$
or
$d_{i}\in\mathscr{E}_{1}=\left\{\alpha_{22},\alpha_{02},\alpha_{20}\right\}\sim\left\{\left(\begin{array}[]{c}2\\ 2\\ 0\end{array}\right),\left(\begin{array}[]{c}0\\ 2\\ 2\end{array}\right),\left(\begin{array}[]{c}2\\ 0\\ 2\end{array}\right)\right\}.$
Suppose $d_{n}\in\mathscr{E}_{1}$ for some large integer $n>n_{0}$, say $d_{n}=\alpha_{22}$. Then by Proposition 2.3 we must have $d_{n-1}=d_{n+1}=\alpha_{11}$. If $d_{n+2}=\alpha_{22}\in\mathscr{E}_{1}$, then we can deduce by using Proposition 2.3 that
$d_{n-1}d_{n}\ldots d_{n+4}\sim\left(\begin{array}[]{c}121211\\ 121211\\ 202022\end{array}\right).$
This implies $\overline{d_{n+3}^{\oplus}d_{n+4}^{\oplus}}=22$, which is forbidden in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},4}$. So we can only have $d_{n+2}\in\left\{\alpha_{21},\alpha_{12}\right\}$. Similar ideas can be applied to deduce that any digit from $\mathscr{E}_{1}$ can not occur in the tail sequence $d_{n+1}d_{n+2}\ldots$. Furthermore, we can apply this idea to prove that there are at most three digits from $\mathscr{E}_{1}$ appearing in the prefix $d_{1}\ldots d_{n}$.
Proof of Lemma 4.5.
@@
(4.17) $d_{m+k}^{1},\penalty 10000\ d_{m+k}^{2},\penalty 10000\ \overline{d_{m+k}^{\oplus}}\leq N+1<M\quad\forall k\geq 0.$
@@
For $d_{m+p+2}$ we have three possibilities: $d_{m+p+2}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{(N+1)(N+1)}\right\}$.
@@
Suppose on the contrary that $d_{m+p+2}=\alpha_{(N+1)(N+1)}$. Then by using $d_{m+p}d_{m+p+1}d_{m+p+2}=\alpha_{(N+1)(N+1)}\alpha_{NN}\alpha_{(N+1)(N+1)}$, and Propositions 2.3 and 4.2 we must have
@@
This implies that $\overline{d_{m+p+2}^{\oplus}}=N-1<M$ and $\overline{d_{m+p+3}^{\oplus}d_{m+p+4}^{\oplus}}=(N+1)(N+1)\succ\delta_{1}(\beta_{c})\delta_{2}(\beta_{c})$, leading to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, Claim 1 holds.
By Claim 1 we have $d_{m+p+2}=\alpha_{(N+1)N}$ or $\alpha_{N(N+1)}$. By symmetry we may assume $d_{m+p+2}=\alpha_{(N+1)N}$. Then by (4.17), and Propositions 2.3 and 4.2 it follows that